INTERMOLECULAR BENDING LEVELS AND INFRARED SPECTROSCOPY OF OPEN-SHELL DIATOM-DIATOM COMPLEXES

\underline{MARK D. MARSHALL},^{a} BETHANY V. POND, SHAWN M. HOPMAN, and MARSHA I. LESTER, Department of Chemistry, University of Pennsylvania, Philadelphia, PA 19104-6323.

A theoretical framework has been developed to describe the bending levels associated with an intermolecular potential of moderate anisotropy between an open-shell diatom and a diatomic partner, such as OH–CO or OH–N_{2}. The model explicitly allows for coupling between the electronic and spin angular momenta of the open-shell OH radical and the vibrational angular momentum arising from the intermolecular bending motions of the complex. Using a dipole-quadrupole interaction potential, energies, wave functions, and transition moments are calculated to reveal the intermolecular energy pattern and expected bending vibrational spectrum for a moderately anisotropic system. The predicted spectrum is compared to that experimentally observed for the OH–N_{2} complex in the OH overtone region. Explicit consideration of the OH radical spin-orbit (l̂· ŝ) and spin-decoupling (ĵ · ŝ) interactions indicates that the spin-orbit splitting in the complex can differ significantly from that of the OH monomer.

aPermanant address: Department of Chemistry, Amherst College, Amherst, MA 01002-5000